### v1

# Every SAM Drop Counts: Embracing Semantic Priors for Multi-Modality Image Fusion and Beyond

Abstract
Multi-modality image fusion, particularly infrared and visible image fusion, plays a crucial role in integrating diverse modalities to enhance scene understanding. Early research primarily focused on visual quality, yet challenges remain in preserving fine details, making it difficult to adapt to subsequent tasks. Recent approaches have shifted towards task-specific design, but struggle to achieve the “The Best of Both Worlds” due to inconsistent optimization goals. To address these issues, we propose a novel method that leverages the semantic knowledge from the Segment Anything Model (SAM) to Grow the quality of fusion results and Establish downstream task adaptability, namely SAGE. Specifically, we design a Semantic Persistent Attention (SPA) Module that efficiently maintains source information via the persistent repository while extracting high-level semantic priors from SAM. More importantly, to eliminate the impractical dependence on SAM during inference, we introduce a bi-level optimization-driven distillation mechanism with triplet losses, which allow the student network to effectively extract knowledge at the feature, pixel, and contrastive semantic levels, thereby removing reliance on the cumbersome SAM model. Extensive experiments show that our method achieves a balance between high-quality visual results and downstream task adaptability while maintaining practical deployment efficiency.
1 Introduction
Figure 1: Differences between the proposed method and existing mainstream comparative approaches: (a) Traditional and early DL-based methods (e.g., U2Fusion) focus on the fusion visual effect. (b) Task-specific methods (e.g., TarDAL & SegMiF) introduce task loss and features that lead to inconsistent optimization goals, causing a conflict between visual and task accuracy. (c) Our pipeline distills the semantic priors of SAM into a student network, achieving practical inference feasibility while ensuring “the best of both worlds” by SAM’s inherent adaptability to these tasks.
In light of recent advancements in sensing technologies, multi-modality imaging has gained significant attention, with applications in robotics [2], remote sensing [26, 25, 14], and autonomous driving [16, 3]. Integrating infrared and visible sensors provides substantial advantages for intelligent processing [18], although both exhibit inherent limitations [24]. Infrared images are robust against smoke, obstruction, and low light, while visible images excel in resolution, contrast, and texture detail. These characteristics are highly complementary, motivating their fusion to generate a comprehensive fused image that balances visual quality and downstream task requirements. Consequently, achieving this fusion efficiently remains a pressing challenge.
Traditional IVIF methods, based on information theory [17], aim to retain as much source image information as possible but often struggle with optimizing fused image quality, especially in managing redundancy and specific scenarios. Early deep learning-based methods [22, 42, 29, 11, 45, 20] focused on visualization but faced issues like edge blurring and artifacts, failing to meet downstream perception needs. Current methods [10, 12, 33, 21] focus on specific tasks, coupling fusion with downstream processing, resulting in conflicting optimization objectives and challenges in balancing visual quality with task adaptability.
Recent advancements in large-scale visual models have significantly improved various visual analysis tasks [38, 5, 40]. Among these, the Segment Anything Model [8] stands out due to its exceptional and robust capability to provide rich semantic information, making it well-suited for IVIF, as illustrated in Figure 2. Also, its segmentation capabilities align naturally with the requirements of downstream tasks in IVIF field. However, current methods that integrate SAM for low-level tasks typically require full SAM during inference, resulting in excessive practical infeasibility.
Figure 2: Comparative robustness of SAM in MFNet normal scenes (top) and challenging FMB conditions (bottom).
To address all the aforementioned issues, we propose a fusion method named SAGE that fully integrates and distills semantic priors derived from SAM. As illustrated in Figure. 1 III, one of our aim is to harness the advantages of SAM’s semantic priors while reducing computational burden. Specifically, the proposed method consists of two key components: Semantic Persistent Attention (SPA) module and a bi-level distillation scheme. The SPA module focuses on using a persistent repository throughout the process to retain source image information, while heuristically enabling the network to integrate semantic patches predicted by SAM. Furthermore, we introduce a distillation scheme designed to efficiently exclude SAM from the inference process, thereby reducing computational complexity. By enhancing distilled information across semantic, feature, and pixel levels, the triplet losses in the scheme significantly boost the performance of the fusion model. Consequently, efficient fusion is achieved using only the distilled model, delivering high-quality visual results and precise task performance without the direct involvement of SAM. In summary, our contributions can be summarized as follows:
  • We utilize semantic priors from SAM to enhance both visual quality and task accuracy for IVIF.
  • A novel Semantic Persistent Attention (SPA) module is proposed, which efficiently integrates semantic priors while preserving source image information by establishing a persistent repository.
  • We develop a multifaceted optimization distillation scheme that transfers the information processed by SPA into the sub-network, effectively decoupling the fusion process from SAM during inference.
  • Extensive experiments on multiple datasets validate the superior effectiveness of our proposed method.
2 Related Works
Learning based IVIF methods. In recent years, deep learning based multi-modality image fusion approaches [22, 42, 29, 11, 45, 20, 50, 49] achieved significant progress, primarily focusing on improving fusion quality, such as enhancing image details, maintaining structural consistency, and reducing noise. These approaches, including innovations in adversarial learning, Transformer models and diffusion network, have provided new insights. However, most of these methods have not addressed how the fused images can directly contribute to downstream tasks such as object detection, semantic segmentation, and other perception-related applications. With the growing demand for more integrated solutions, this gap has sparked increasing research interest in combining fusion with various downstream tasks.
Early works bridged this gap, with TarDAL [10] introducing a dual-discriminator network for detection optimization, jointly addressing modality differences and detection tasks. Recent methods [36, 37] further focused on detection-optimized fusion. For semantic segmentation [44], SeAFusion [33] and SegMiF [12] improved fusion through task loss and task-specific features, respectively. Recent approaches [13, 21, 32, 19] integrated multi-task learning to jointly optimize detection and segmentation, enhancing performance for both tasks.
Figure 3: A overall workflow of our proposed method. (a) shows the flow structure of the main network, where the SPA module processes patches with semantic priors generated by SAM. (b) illustrates the detailed structure of the SPA module, where PR plays a key role in preserving the source and integrating the semantic information. (c) displays our distillation scheme formulation, with visualizations of the different components of the triplet loss. (d) provides a simple diagram of the sub-network, which is composed of stacked Dense blocks.
SAM & its application.  The Segment Anything Model (SAM) [8] is a large-scale pre-trained foundation model [38] that excels in zero-shot generalization, enabling it to handle a wide range of segmentation and detection tasks without the need for task-specific training. SAM demonstrates impressive performance in low-level vision applications, such as image deblurring [9], dehazing [15], super-resolution [23, 35], and low-light image enhancement [6]. Moreover, SAM’s capabilities extend to high-level tasks like object detection [30], which further strengthens its versatility. This ability to handle both low-level and high-level tasks makes SAM inherently well-suited for the needs of the IVIF field, where fusion based segmentation and detection tasks are critical. However, despite its impressive performance, the high computational complexity presents significant challenges for its practical deployment. Addressing these computational burdens, while fully leveraging semantic priors of SAM, is one of the main primary motivations for our work.
3 The Proposed Method
3.1 Motivation and Problem Formulation
Figure 4: Visual comparison of SOTA approaches across four commonly used datasets, including TNO, RoadScene, M3FD and FMB.
The overall workflow of the proposed method is shown in Figure 3. Our objective is to develop a framework that effectively learns and leverages the semantic priors provided by SAM while minimizing the computational burden during inference. Integrating SAM’s semantic information can significantly enhance the quality of the fused images; however, directly involving SAM make the inference process computationally impractical due to its complexity. To solve this, we utilize a distillation scheme to transfer knowledge from a SAM-integrated main network to a lightweight sub-network. Traditional distillation approaches often face difficulties in capturing the intricate semantic relationships and maintaining the structural consistency necessary for tasks involving cross-modal information fusion. Considering these challenges and the need to jointly optimize both networks for effective semantic transfer, we propose a bi-level optimization framework that models the main and sub-networks as a unified system, aiming to enable a more cohesive and structured transfer of semantic insights.
Denote that $\textbf{I}_{\text{vis}}$ and $\textbf{I}_{\text{ir}}$ represent the visible and infrared input images, respectively. Their corresponding semantic prior patches generated from SAM can be defined as $\textbf{P}_{\text{vis}}$ and $\textbf{P}_{\text{ir}}$. The main network, represented by $\mathcal{M}$ with parameters $\boldsymbol{\omega_{\mathrm{m}}}$. while the sub-network, represented by $\mathcal{S}$ with parameters $\boldsymbol{\omega_{\mathrm{s}}}$, aims to distill knowledge from the main network through the distillation process. Below our definition, the optimization process can be formulated as follows:
$\min_{\omega_{\text{s}}}\mathcal{L}^{\text{d}}_{\text{m}}\big{(}\mathcal{M}(\{\textbf{I}_{\text{vis}},\textbf{I}_{\text{ir}}\};\{\textbf{P}_{\text{vis}},\textbf{P}_{\text{ir}}\};\boldsymbol{\omega_{\mathrm{m}}})\big{)},$ (1)
$s.t.\boldsymbol{\omega_{\mathrm{s}}}^{*}\in\arg\min\limits_{\boldsymbol{\omega_{\mathrm{s}}}}\mathcal{L}^{\mathrm{d}}_{\text{s}}\big{(}\mathcal{S}(\mathbf{I_{\mathrm{vis}}},\mathbf{I_{\mathrm{ir}}};\boldsymbol{\omega_{\mathrm{s}}})\big{)},$ (2)
where, $\mathcal{L}^{\mathrm{d}}_{\text{m}}$ represents the distillation loss for main network, ensuring high-quality fusion while embedding meaningful semantic information, and $\mathcal{L}^{\mathrm{d}}_{\text{s}}$ represents the distillation loss to train the sub-network for effectively mimicking the behavior of the main network. The optimized interaction process is reflected in fused refer image from $\mathcal{M}$, $\mathbf{I_{\mathrm{ref}}}=\mathcal{M}(\{\textbf{I}_{\text{vis}},\textbf{I}_{\text{ir}}\};\{\textbf{P}_{\text{vis}},\textbf{P}_{\text{ir}}\},\omega_{\text{m}})$ and the fused image $\mathbf{I_{\mathrm{f}}}=\mathcal{S}(\mathbf{I_{\mathrm{vis}}},\mathbf{I_{\mathrm{ir}}};\boldsymbol{\omega_{\mathrm{s}}}^{*})$. The goal is to jointly optimize both networks, where the main network learns the rich semantic priors, and the sub-network captures those features under reduced computational requirements.
In the main network, source images are processed by SAM to generate image patches with semantic priors, which, along with features extracted from the source images at multiple scales, are passed through the SPA module to produce fused images. The sub-network serves as a distilled version of the main network. The bi-level optimization distillation scheme, involving triplet losses, facilitates the structured transfer of capabilities from the main network to the sub-network, ensuring that even with reduced complexity. The following sections will detail the design of the SPA module (Section 3.2) and the bi-level optimization distillation scheme involving triplet losses (Section 3.3).
3.2 Semantic Persistent Attention Module
To fully exploit the semantic information provided by SAM, we propose the Semantic Persistent Attention (SPA) Module, as illustrated in the flowchart in Figure. 3 (b).
The key core of the SPA module is a Persistent Repository (PR), which serves as a static memory to store and maintain crucial contextual information for the fusion process. Specifically, PR stores the latent representation (Z) of the source features $\textbf{F}_{\text{src}}$ and the corresponding key-value pairs ($\textbf{K}_{\text{src}},\textbf{V}_{\text{src}}$), which provide consistent contextual support during the cross-attention operation. The core insight behind this design is that PR acts as a stable, modality-specific information source, guiding the cross-attention mechanism to fuse the semantic patches with rich contextual details from the original source.
The visible and infrared semantic patches ($\textbf{P}_{\text{vis}}$ and $\textbf{P}_{\text{ir}}$) extracted from the Semantic Attention Module (SAM) are encoded into features ($\textbf{F}_{\text{pvis}}$ and $\textbf{F}_{\text{pir}}$). These encoded semantic features represent limited portions of the scene and are processed through the cross-attention mechanism. By utilizing the key-value pairs ($\textbf{K}_{\text{src}},\textbf{V}_{\text{src}}$) stored in PR, the cross-attention mechanism enriches the semantic queries ($\textbf{Q}_{\text{pvis}}$ and $\textbf{Q}_{\text{pir}}$) with the full scene context from the source image, thereby addressing the inherent incomplete scene coverage in the patches. PR ensures that the fusion process remains grounded in a stable and consistent context, allowing the patches to be effectively supplemented with modality-specific information from the source. This stable context prevents the fusion process from drifting or losing semantic coherence, while still enabling flexibility in refining the features through attention mechanisms.
In summary, the SPA module uses PR to guide the cross-attention mechanism, ensuring that the semantic patches from SAM are enriched with consistent and detailed modality-specific information. This approach allows the fusion of infrared and visible image modalities by combining both the high-level semantic understanding from SAM and the detailed information from the source features. The final output, $\textbf{F}_{\text{SPA}}$, represents a semantically enriched, structurally consistent feature set, rich in semantic priors of SAM, and ready for further processing.
The main network heavily utilizes the SPA module, aiming to fully exploit the complex semantic priors provided by SAM, focus on capturing and retaining detailed semantic knowledge, thereby facilitating a powerful representation that can later be distilled into a more efficient sub-network.
Figure 5: Quantitative comparison of fusion performance with other state-of-the-art methods on M3FD and FMB benchmarks. The Violin plots illustrating the distribution of metrics, in which the the black lines and white triangles indicate the medium and mean values.
3.3 Triplet Loss in Distillation Scheme
In this section, we introduce a triplet loss-driven distillation scheme that facilitates semantic knowledge transfer from the main network to the sub-network, where both networks are optimized in a bidirectional manner. This allows gradients to flow both from the main to the sub-network and vice versa, ensuring a cohesive transfer with mutual compromise. The distillation is driven by three distinct loss components, each targeting different aspects.
The first loss operates at the feature level, aligning the reference image ($\mathbf{I}_{\text{ref}}$) and the fused image ($\mathbf{I}_{\text{f}}$) across multiple scales, which is defined as:
$\mathcal{L}_{\text{fea}}=\sum_{m=1}^{M}\left(1-\frac{\textbf{F}_{\text{Den}}^{m}\cdot\textbf{F}_{\text{SPA}}^{m}}{\|\textbf{F}_{\text{Den}}^{m}\|_{2}\cdot\|\textbf{F}_{\text{SPA}}^{m}\|_{2}}\right),$ (3)
where $\textbf{F}_{\text{Den}}^{m}$ and $\textbf{F}_{\text{SPA}}^{m}$ are the feature maps of the Dense and SPA blocks at the same scale $m$, and $\|\cdot\|_{2}$ is the l2-norm.
The second loss component operates at the context level, which consists of two parts: a gradient part to preserve structural consistency and an MSE loss for intensity consistency. They are defined as:
$\mathcal{L}_{\text{grad}}=\left\|\nabla\mathbf{I}_{\text{ref}}-\nabla\mathbf{I}_{\text{fus}}\right\|_{1},\quad\mathcal{L}_{\text{MSE}}=\left\|\mathbf{I}_{\text{ref}}-\mathbf{I}_{\text{fus}}\right\|_{2},$ (4)
where $\nabla$ denotes the Sobel operator and $\|\cdot\|_{1}$ is the l1-norm. So the cont loss can be calculated as $\mathcal{L}_{\text{context}}=\mathcal{L}_{\text{grad}}+\mathcal{L}_{\text{MSE}}$.
The third loss component, contrastive semantic loss aims to ensure that the feature space of the fused image remains close to that of the reference image, while being distinct from the individual visible and infrared images archiving efficient distillation. We utilize the semantic feature space by encoder of SAM (defined as $\mathcal{S}_{\text{E}}$) to construct positive and negative pairs through the element-wise multiplication of the binary masks ($\mathbf{M}_{\text{vis}}$ and $\mathbf{M}_{\text{ir}}$) from $\textbf{P}_{\text{vis}}$ and $\textbf{P}_{\text{ir}}$. The contrastive semantic loss on each modality is defined as:
$\mathcal{L}_{\text{cs}}^{\text{ir}}=\sum_{l=1}^{L}\frac{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{fus}}\odot\mathbf{M}_{\text{ir}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ref}}\odot\mathbf{M}_{\text{ir}})\right\|_{1}^{2}}{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{x}}\odot\mathbf{M}_{\text{ir}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ir}}\odot\mathbf{M}_{\text{ir}})\right\|_{1}^{2}},$ (5)
$\mathcal{L}_{\text{cs}}^{\text{vis}}=\sum_{l=1}^{L}\frac{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{fus}}\odot\mathbf{M}_{\text{vis}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ref}}\odot\mathbf{M}_{\text{vis}})\right\|_{1}^{2}}{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{x}}\odot\mathbf{M}_{\text{vis}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{vis}}\odot\mathbf{M}_{\text{vis}})\right\|_{1}^{2}},$ (6)
where $l$ means $l$th layer of $\mathcal{S}_{\text{E}}$, and x $\in\{\text{ref},\text{fus}\}$. So the total $\mathcal{L}_{\text{cs}}=\mathcal{L}_{\text{cs}}^{\text{ir}}+\mathcal{L}_{\text{cs}}^{\text{vis}}$ can also be computed bidirectionally in the bi-level optimization scheme between two networks, ensuring that the feature spaces of both the $\mathbf{I}_{\text{ref}}$ and the $\mathbf{I}_{\text{f}}$ are easily aligned.
The total distillation loss for the sub-network, denoted as $\mathcal{L}^{\text{d}}_{\text{s}}$ in Eq. 2, is the sum of the feature loss, context loss, and contrastive semantic loss: $\mathcal{L}^{\text{d}}_{\text{s}}=\mathcal{L}_{\text{fea}}+\mathcal{L}_{\text{cont}}+\mathcal{L}_{\text{cs}}$.
For the main network, an additional segmentation cross-entropy loss $\mathcal{L}^{\text{seg}}$ is incorporated. This loss is computed between the segmentation predictions generated by an open-vocabulary SAM model $\mathcal{S}^{\text{O}}$ and the ground truth segmentation labels. By including this loss, we aim to leverage the capabilities of SAM to overcome the limitations of traditional segmentation networks, which typically rely on a fixed vocabulary. Instead of affecting the sub-network, this loss aim to prevent potential optimization conflicts. The segmentation loss $\mathcal{L}^{\text{seg}}$ is defined as:
$\mathcal{L}^{\text{seg}}=-\sum_{c}\left[\mathbf{Y}_{\text{label}}^{c}\log(\mathbf{Y}_{\text{pred}}^{c})\right],$ (7)
where $\mathbf{Y}_{\text{label}}$ represents the label segmentation map, $\mathbf{Y}_{\text{pred}}$ is the predicted segmentation map, and $c$ is the class index. For now, the total distillation loss for the main network can be calculated as: $\mathcal{L}^{\text{d}}_{\text{m}}=\mathcal{L}^{\text{d}}_{\text{s}}+\mathcal{L}^{\text{seg}}$.
4 Experiments
Framework Segformer - B3 [41] X - Decoder [51]
Method Venue Road Bldg. Sign Veg. Bus Per. Car mIoU Road Bldg. Sign Veg. Bus Per. Car mIoU
Visible - 84.8 81.0 64.2 84.7 46.8 61.1 77.1 49.1 79.2 67.8 18.2 73.2 52.0 56.9 82.0 50.1
Infrared - 83.7 77.6 55.9 80.9 38.1 57.1 74.2 42.7 76.6 55.6 12.9 38.2 72.7 63.2 78.7 42.9
DDFM ICCV’23 86.2 83.0 69.5 84.8 62.5 61.0 80.8 58.2 76.1 65.9 10.1 73.3 72.0 66.6 82.5 49.3
U2Fusion TPAMI’22 86.7 82.4 68.5 84.6 50.3 61.8 80.4 57.0 79.7 70.6 22.2 72.7 75.7 65.5 83.3 50.4
TarDAL CVPR’22 85.8 81.8 68.4 84.7 49.5 61.3 79.5 55.7 74.1 63.1 23.2 67.5 74.8 66.4 81.1 48.7
SegMiF ICCV’23 86.3 82.2 68.0 85.0 63.7 60.7 80.0 57.3 79.6 68.6 14.0 73.8 74.3 65.8 83.2 50.8
FILM ICML’24 86.9 81.9 69.1 84.0 46.1 60.6 79.4 55.7 79.2 70.3 17.9 73.7 79.0 64.0 82.0 50.6
SHIP CVPR’24 84.9 82.1 68.3 83.3 61.4 60.9 80.4 57.2 76.4 66.8 24.7 71.2 76.3 64.6 80.9 50.4
MRFS CVPR’24 88.2 82.8 71.0 86.0 61.5 63.6 81.1 56.5 73.8 65.0 23.7 72.3 77.2 66.2 82.6 49.6
EMMA CVPR’24 85.2 81.2 66.8 83.3 59.0 58.9 79.9 55.8 78.6 68.4 15.0 74.3 77.1 64.5 82.2 50.7
TIMFusion TPAMI’24 85.1 80.0 66.4 82.8 50.8 59.0 79.7 55.5 62.0 63.0 12.8 67.6 75.8 65.8 79.6 46.1
SAGE Proposed 89.0 83.9 71.4 86.3 66.7 64.1 81.7 61.2 81.0 70.4 17.0 75.2 77.7 64.0 82.8 51.1
Table 1: Quantitative semantic segmentation IoU results of different methods on the FMB dataset.
Figure 6: Visual comparison of segmentation results of different methods on the FMB dataset.
4.1 Settings and Details
Five representative datasets, namely TNO [34], RoadScene [42], MFNet [4], FMB [12], and M3FD [10], are utilized for both training and evaluation. For fusion-based segmentation evaluations, SegFormer [41] (B3 variant) is adopted as the backbone, and the models are trained for 100 epochs. The training and testing splits are strictly in accordance with the official dataset guidelines. The Adam optimizer is used to optimize both networks during the training period, with learning rates of $5\times 10^{-5}$ for both networks. Prior to distillation, the models undergo a pre-training phase, followed by 5 epochs of distillation. The batch size is set to 4, and the images are randomly cropped and resized to $192\times 160$ during training. The entire framework is implemented in PyTorch and executed on two NVIDIA GeForce RTX 4090 GPUs.
4.2 Results of IVIF
We demonstrate our fusion quality by visualization and quantitative analysis with 9 other SOTA methods in recent years, including DDFM [45], U2Fusion [42], TarDAL [10], SegMiF [12], FILM [47], SHIP [48], MRFS [44], EMMA [46] and TIMFusion [19].
Visualization Comparisons. Figure. 4 presents a visual comparison of various methods. Overall, the proposed method has two main advantages. First, it effectively preserves the multi-modality information of the original images. In the TNO surveillance scene, the vegetation details from the visible image and the smoke from the chimney in the infrared image are both well retained (the image set in the top-left corner). In RoadScene, our method also achieves optimal leaf recovery. On the other hand, our approach demonstrates strong robustness to interference, as it accurately reconstructs the reflective pedestrian crossing lines on the ground and the distant building outlines in the dense fog at night (the green box in the second row). The integration of SAM enhances our method, giving it the ability to surpass other methods and achieve superior results.
Quantitative Results. We also compare our numerical results with other competitors in Figure. 5, based on 100 image pairs randomly selected from the FMB and M3FD datasets. We employ four fusion evaluation metrics, including the Entropy (EN) [31], Standard Deviation (SD) [1], Sum of Correlation Differences (SCD) [24], and Multi-Scale Structural Similarity (MS-SSIM) [39]. Additionally, we introduce four No-Reference Image Quality Assessments (NR-IQA) related to image quality: BRISQUE [27], NIQE [28], MUSIQ [7], and PaQ-2-PiQ [43].
In terms of these metrics, our method demonstrates consistent superiority. High SCD and MS-SSIM values indicate that our method effectively preserves a significant amount of information from the source images, while EN reflects our rich detail and texture. The spatial, multi-scale, and patch evaluations of the NR-IQA confirm that our fusion approach has a better pixel distribution and aligns more closely with the human visual system perception.
Framework Segformer - B3 [41]
Method Venue Car Per. Bike Curve C.S. C.C. Bump mIoU
Visible - 83.3 54.7 59.6 7.46 31.0 37.4 30.8 44.6
Infrared - 79.0 63.8 51.4 7.10 25.2 34.1 41.8 44.4
DDFM ICCV’23 86.7 62.2 63.4 30.8 35.5 45.7 41.1 51.7
U2Fusion TPAMI’22 87.3 65.9 60.8 33.4 36.0 46.1 40.9 52.0
TarDAL CVPR’22 86.4 67.1 62.4 28.4 30.5 45.0 41.5 51.0
SegMiF ICCV’23 85.9 66.3 64.4 34.0 34.7 43.0 48.0 52.4
FILM ICML’24 86.6 66.4 61.9 36.3 32.0 46.7 46.7 52.7
SHIP CVPR’24 86.7 63.9 59.7 33.5 35.2 44.4 42.4 51.5
MRFS CVPR’24 86.1 61.5 59.7 30.6 35.2 44.9 38.4 50.4
EMMA CVPR’24 86.1 65.6 63.8 25.4 29.1 44.9 41.3 51.7
TIMFusion TPAMI’24 86.4 58.7 60.3 29.8 33.6 44.1 38.3 49.8
SAGE Proposed 87.9 67.5 64.3 41.2 34.1 46.8 46.7 54.0
Table 2: Quantitative semantic segmentation results of different methods on the MFNet dataset.
Figure 7: Visual comparison of segmentation results of different methods on the MFNet dataset.
4.3 Results of IVIF based Segmentation
Visualization Comparisons. We present the segmentation visualization results on the challenging new FMB dataset in Figure. 6. In the daytime intersection scene, thanks to the powerful semantic priors provided by SAM, our method is the only one to completely distinguish between the truck and the bus. In the nighttime road scene, we successfully segment the sidewalk, achieving optimal performance. Additionally, we showcase the segmentation results on the MFNet dataset in Figure. 7. Leveraging the advanced semantic information learned by SAM, our method successfully segments distant small pedestrian targets during the daytime and near-invisible lane curve at night.
Quantitative Results. Table. 1 presents a qualitative comparison of IoU results for two segmentation frameworks on the FMB dataset. The first framework involves re-training images generated by various fusion methods, while the second uses the output from the open-vocabulary segmentation network with prompt words, without requiring retraining. In traditional comparisons, our method achieves a 3.0 mIoU improvement over the second-best approach, demonstrating its competitive edge across multiple categories. Additionally, in the segmentation network that requires no training, our method also performs favorably, benefiting from the efficient adaptability provided by SAM’s semantic information. Table. 2 further showcases the qualitative results on the MFNet dataset, which has lower resolution and fewer labels. Although our method shows some competitive gaps in background categories (e.g., bicycles, colored cones), it achieves significant overall superiority in key perception categories (e.g., Car and Person) and mean performance. Our method seamlessly integrates SAM with the fusion network, achieving a breakthrough in performance for efficient segmentation task.
Variants mIoU Metric
MFNet / FMB EN / SD/ SCD / MS-SSIM
I. SAM in the Pipeline
(a) w/o SAM in $\mathcal{M}$ 50.7 / 56.5 6.713 / 36.93 / 1.612 / 1.127
(b) P from $\mathbf{Y}_{\text{label}}$ 53.2 / 56.4 6.239 / 33.52 / 1.412 / 1.102
(c) Swap $\mathcal{S}^{\text{O}}$ 53.4 / 55.5 6.312 / 34.79 / 1.316 / 1.097
II. Design of SPA
(a) w/o Z 52.3 / 57.2 6.333 / 32.81 / 1.384 / 1.002
(b) w/o k-v in PR 52.1 / 56.1 6.417 / 37.19 / 1.261 / 1.063
(c) w/o PR 53.3 / 57.2 6.556 / 35.48 / 1.576 / 1.046
III. Distillation Triplet Losses
(a) w/o $\mathcal{L}_{\text{fea}}$ 49.7 / 51.6 6.436 / 37.28 / 1.277 / 1.079
(b) w/o $\mathcal{L}_{\text{cont}}$ 50.9 / 56.2 6.669 / 37.80 / 1.462 / 1.086
(c) w/o $\mathcal{L}_{\text{cs}}$ 51.7 / 54.3 6.524 / 39.11 / 1.356 / 1.038
(d) Online Dist. 50.0 / 50.4 6.221 / 32.65 / 1.581 / 0.987
The proposed method 54.0 / 61.2 6.872 / 40.77 / 1.604 / 1.117
Table 3: Quantitative results on both fusion and segmentation of all variants in three ablation studies.
Method DDFM U2Fusion TarDAL SegMiF FILM SHIP MRFS EMMA TIMFusion Ours
Time(ms) 280K 42.31 16.57 150.5 183.3 27.93 95.86 25.73 17.39 10.47
FLOPS(G) 1.34M 518.0 233.3 500.4 230.3 401.5 301.2 106.4 100.7 52.06
Params(M) 552.7 0.659 0.297 0.621 0.491 0.525 133.4 1.516 0.158 0.136
Table 4: Efficiency comparison with other state-of-the-art methods on M3FD benchmarks.
4.4 Ablation Studies
Impact of SAM. The Segment Anything Module serves as a critical cornerstone in our method. To explore its impact, we derive three variants by isolating SAM from the core framework. Specifically, in variant (a), we replace the semantic patches with randomly cropped source image patches, thereby removing SAM from the main network $\mathcal{M}$. In variant (b), we assist the generation of semantic patches with segmentation labels. Additionally, we replace $\mathcal{S}^{\text{O}}$ with a conventional segmentation network [41]. Variant (a) enhances the source image information, leading to the optimal similarity metrics, SCD and MS-SSIM, as shown in Table. 3. Our method fully leverages the semantic priors provided by SAM, resulting in the best foggy building contours, as illustrated in the first row of Figure. 8.
Figure 8: Visualization comparison of impact of SAM.
Figure 9: Visualization comparison of ablation on SPA.
Study on SPA module. The proposed Semantic Persistent Attention module plays a key role in integrating the semantic priors from SAM, effectively guiding the retention and enhancement of inherent modality features. To investigate its impact, we set up three variants: (a) w/o latent representation (Z), (b) w/o key-value pairs in the Persistent Repository (PR), and (c) w/o PR. The channel loss due to these modifications has been corrected. A visual comparison of these ablation variants is shown in Figure. 9. It is evident that PR in the SPA module plays an essential role in maintaining critical information. The removal of any component leads to significant information loss. For instance, in (b), while the regions of interest are highlighted, the absence of source image information causes blurriness. Similarly, the absence of PR in the SPA module completely prevents the capture of beneficial semantic information, causing the network to lose focus on key areas and resulting in a low-contrast fusion results in (c).
Discussion on distillation scheme. Furthermore, we discuss the distillation scheme. We first conduct ablation studies on each component of the triple loss function, yielding variants (a)-(c) in Table. 3 III. We also replace the dual-layer optimization distillation with a traditional online distillation approach, forming variant (d). A visual comparison is shown in Figure. 10. Notably, traditional online distillation results are significantly inferior to our method, both in visual quality and the gradient map from variant (b). This confirms the effectiveness of our distillation method when incorporating semantic information.
Figure 10: Visualization comparison of ablation on Distillation.
4.5 Computational Efficiency Analysis
In the M3FD benchmark, we compare our method with 9 other SOTA methods in terms of time, FLOPS, and parameters. As shown in Table. 4, our method demonstrates significant advantages across all aspects, particularly in time and FLOPS. Specifically, due to our distillation scheme, our sub-network is able to adjust flexibly while maintaining high efficiency, resulting in reduced FLOPS. Compared to other methods, our approach achieves a processing time of 10.47 ms and FLOPS of 52.06 G, outperforming most of the existing methods, with a parameter count of only 0.136M, showcasing computational efficiency.
Our method effectively alleviates the computational burden of SAM during inference, significantly reducing the computational overhead while preserving semantic information. This design allows our network to strike a balance between processing speed and computational resources, leading to superior efficiency in practical applications.
5 Conclusion
In this work, we propose a fusion method that utilizes semantic priors from the SAM for IVIF. We design the Semantic Persistent Attention module, which integrates semantic information while retaining source details. Additionally, we introduce a bi-level distillation scheme with triplet loss to reduce computational complexity by decoupling fusion from SAM during inference. Extensive experiments show that our method achieves superior fusion and task performance, outperforming across multiple datasets.

### v2

@@
 Abstract
-Multi-modality image fusion, particularly infrared and visible image fusion, plays a crucial role in integrating diverse modalities to enhance scene understanding.
-Early research primarily focused on visual quality, yet challenges remain in preserving fine details, making it difficult to adapt to subsequent tasks.
-Recent approaches have shifted towards task-specific design, but struggle to achieve the “The Best of Both Worlds” due to inconsistent optimization goals.
-To address these issues, we propose a novel method that leverages the semantic knowledge from the Segment Anything Model (SAM) to Grow the quality of fusion results and Establish downstream task adaptability, namely SAGE.
+Multi-modality image fusion, particularly infrared and visible, plays a crucial role in integrating diverse modalities to enhance scene understanding.
+Although early research prioritized visual quality, preserving fine details and adapting to downstream tasks remains challenging.
+Recent approaches attempt task-specific design but rarely achieve “The Best of Both Worlds” due to inconsistent optimization goals.
+To address these issues, we propose a novel method that leverages the semantic knowledge from the Segment Anything Model (SAM) to Grow the quality of fusion results and Enable downstream task adaptability, namely SAGE.
 Specifically, we design a Semantic Persistent Attention (SPA) Module that efficiently maintains source information via the persistent repository while extracting high-level semantic priors from SAM.
-More importantly, to eliminate the impractical dependence on SAM during inference, we introduce a bi-level optimization-driven distillation mechanism with triplet losses, which allow the student network to effectively extract knowledge at the feature, pixel, and contrastive semantic levels, thereby removing reliance on the cumbersome SAM model.
+More importantly, to eliminate the impractical dependence on SAM during inference, we introduce a bi-level optimization-driven distillation mechanism with triplet losses, which allow the student network to effectively extract knowledge.
 Extensive experiments show that our method achieves a balance between high-quality visual results and downstream task adaptability while maintaining practical deployment efficiency.
+The code is available at https://github.com/RollingPlain/SAGE_IVIF.
 1 Introduction
-Figure 1: Differences between the proposed method and existing mainstream comparative approaches: (a) Traditional and early DL-based methods (e.g., U2Fusion) focus on the fusion visual effect.
-(b) Task-specific methods (e.g., TarDAL & SegMiF) introduce task loss and features that lead to inconsistent optimization goals, causing a conflict between visual and task accuracy.
-(c) Our pipeline distills the semantic priors of SAM into a student network, achieving practical inference feasibility while ensuring “the best of both worlds” by SAM’s inherent adaptability to these tasks.
-In light of recent advancements in sensing technologies, multi-modality imaging has gained significant attention, with applications in robotics [2], remote sensing [26, 25, 14], and autonomous driving [16, 3].
-Integrating infrared and visible sensors provides substantial advantages for intelligent processing [18], although both exhibit inherent limitations [24].
+Figure 1: Differences between the proposed method and existing mainstream comparative approaches: (a) Traditional and early DL-based methods focus on the fusion visual effect.
+(b) Task-specific methods (e.g., TarDAL [11] & SegMiF [13]) introduce task loss and features that lead to inconsistent optimization goals, causing a conflict between visual and task accuracy.
+(c) Our pipeline first leverages semantic priors from SAM within a large network and then distills the knowledge into a smaller sub-network achieving practical inference feasibility while ensuring “the best of both worlds” through SAM’s inherent adaptability to these tasks.
+In light of recent advancements in sensing technologies, multi-modality imaging [15] has gained significant attention, with applications in robotics [35], remote sensing [33], and autonomous driving [1].
+Integrating infrared and visible sensors provides substantial advantages for intelligent processing [5], although both exhibit inherent limitations [21, 16].
 Infrared images are robust against smoke, obstruction, and low light, while visible images excel in resolution, contrast, and texture detail.
 These characteristics are highly complementary, motivating their fusion to generate a comprehensive fused image that balances visual quality and downstream task requirements.
-Consequently, achieving this fusion efficiently remains a pressing challenge.
-Traditional IVIF methods, based on information theory [17], aim to retain as much source image information as possible but often struggle with optimizing fused image quality, especially in managing redundancy and specific scenarios.
-Early deep learning-based methods [22, 42, 29, 11, 45, 20] focused on visualization but faced issues like edge blurring and artifacts, failing to meet downstream perception needs.
+Consequently, achieving this fusion efficiently remains a pressing challenge [15].
+Traditional Infrared and Visible Image Fusion (IVIF) methods, based on information theory [37, 17], aim to retain as much source image information as possible but often struggle with optimizing fused image quality, especially in managing redundancy and specific scenarios.
+Early deep learning-based methods [32, 8, 12, 40, 39, 22, 18] focused on visualization but faced issues like edge blurring and artifacts, failing to meet downstream perception needs.
 Current methods [11, 13, 26, 20] focus on specific tasks, coupling fusion with downstream processing, resulting in conflicting optimization objectives and challenges in balancing visual quality with task adaptability.
-Recent advancements in large-scale visual models have significantly improved various visual analysis tasks [38, 5, 40].
+Recent advancements in large-scale visual models have significantly improved various visual analysis tasks [30].
 Among these, the Segment Anything Model [6] stands out due to its exceptional and robust capability to provide rich semantic information, making it well-suited for IVIF, as illustrated in Figure 2.
@@
 However, current methods that integrate SAM for low-level tasks typically require full SAM during inference, resulting in excessive practical infeasibility.
-Figure 2: Comparative robustness of SAM in MFNet normal scenes (top) and challenging FMB conditions (bottom).
-To address all the aforementioned issues, we propose a fusion method named SAGE that fully integrates and distills semantic priors derived from SAM.
-As illustrated in Figure.
-1 III, one of our aim is to harness the advantages of SAM’s semantic priors while reducing computational burden.
-Specifically, the proposed method consists of two key components: Semantic Persistent Attention (SPA) module and a bi-level distillation scheme.
-The SPA module focuses on using a persistent repository throughout the process to retain source image information, while heuristically enabling the network to integrate semantic patches predicted by SAM.
-Furthermore, we introduce a distillation scheme designed to efficiently exclude SAM from the inference process, thereby reducing computational complexity.
-By enhancing distilled information across semantic, feature, and pixel levels, the triplet losses in the scheme significantly boost the performance of the fusion model.
+Figure 2: Demonstration of SAM’s robustness in MFNet normal scenes (top) and under challenging FMB conditions (bottom).
+To address these issues, we propose the fusion method SAGE that fully integrates and distills semantic priors derived from SAM.
+As illustrated in Figure 1 III, one of our aim is to harness the advantages of SAM’s semantic priors while reducing computational burden.
+The proposed method consists of two key components: Semantic Persistent Attention module and a bi-level distillation scheme.
+The former focuses on using a persistent repository throughout the process to retain source image information, while heuristically enabling the network to integrate semantic patches predicted by SAM.
+Furthermore, the distillation scheme is designed to efficiently exclude SAM from the inference process, thereby reducing computational complexity.
+By enhancing distilled information across different aspects, the triplet losses in the scheme significantly boost the performance of the fusion model.
 Consequently, efficient fusion is achieved using only the distilled model, delivering high-quality visual results and precise task performance without the direct involvement of SAM.
 In summary, our contributions can be summarized as follows:
-  • We utilize semantic priors from SAM to enhance both visual quality and task accuracy for IVIF.
-  • A novel Semantic Persistent Attention (SPA) module is proposed, which efficiently integrates semantic priors while preserving source image information by establishing a persistent repository.
-  • We develop a multifaceted optimization distillation scheme that transfers the information processed by SPA into the sub-network, effectively decoupling the fusion process from SAM during inference.
+  • A novel fusion framework is proposed to leverage semantic priors from the SAM, effectively balancing visual quality and downstream task adaptability..
+  • The designed Semantic Persistent Attention (SPA) module utilizes a persistent repository to efficiently retain source information and simultaneously extract high-level semantic representation.
+  • We develop a bi-level optimization distillation scheme that transfers the information processed by SPA into the sub-network, effectively decoupling the fusion process from SAM during inference.
   • Extensive experiments on multiple datasets validate the superior effectiveness of our proposed method.
@@
 Learning based IVIF Methods.
-In recent years, deep learning based multi-modality image fusion approaches [22, 42, 29, 11, 45, 20, 50, 49] achieved significant progress, primarily focusing on improving fusion quality, such as enhancing image details, maintaining structural consistency, and reducing noise.
+In recent years, deep learning based multi-modality image fusion approaches [32, 8, 12, 40, 39, 44] achieved significant progress, primarily focusing on improving fusion quality, such as enhancing image details, maintaining structural consistency, and reducing noise.
 These approaches, including innovations in adversarial learning, Transformer models and diffusion network, have provided new insights.
@@
 With the growing demand for more integrated solutions, this gap has sparked increasing research interest in combining fusion with various downstream tasks.
-Early works bridged this gap, with TarDAL [10] introducing a dual-discriminator network for detection optimization, jointly addressing modality differences and detection tasks.
-Recent methods [36, 37] further focused on detection-optimized fusion.
-For semantic segmentation [44], SeAFusion [33] and SegMiF [12] improved fusion through task loss and task-specific features, respectively.
-Recent approaches [13, 21, 32, 19] integrated multi-task learning to jointly optimize detection and segmentation, enhancing performance for both tasks.
+Early works began to bridge this gap, with TarDAL [11] pioneering the use of a dual-discriminator network for object detection optimization, which simultaneously considers modality differences and detection tasks during fusion.
+As the field has progressed, several methods [24, 38] have also focused on optimizing fusion for object detection.
+For semantic segmentation [36], SeAFusion [26] and SegMiF [13] have made early significant improvements, with the former enhancing fusion via task loss and the latter leveraging task-specific features.
+More recently, some approaches [14, 20, 25, 19, 3, 29] have integrated multi-task learning to jointly optimize detection and segmentation, improving performance on both tasks.
 Figure 3: A overall workflow of our proposed method.
@@
 SAM & Its Application.
-The Segment Anything Model (SAM) [8] is a large-scale pre-trained foundation model [38] that excels in zero-shot generalization, enabling it to handle a wide range of segmentation and detection tasks without the need for task-specific training.
-SAM demonstrates impressive performance in low-level vision applications, such as image deblurring [9], dehazing [15], super-resolution [23, 35], and low-light image enhancement [6].
+The Segment Anything Model (SAM) [6] is a large-scale pre-trained foundation model that excels in zero-shot generalization, enabling it to handle a wide range of segmentation and detection tasks without the need for task-specific training.
+SAM demonstrates impressive performance in low-level vision applications, such as image deblurring [9], dehazing [4], super-resolution [28], and low-light image enhancement [7].
 Moreover, SAM’s capabilities extend to high-level tasks like object detection [23], which further strengthens its versatility.
 This ability to handle both low-level and high-level tasks makes SAM inherently well-suited for the needs of the IVIF field, where fusion based segmentation and detection tasks are critical.
-However, despite its impressive performance, the high computational complexity presents significant challenges for its practical deployment.
-Addressing these computational burdens, while fully leveraging semantic priors of SAM, is one of the main primary motivations for our work.
+However, its impressive performance is hindered by high computational complexity, posing challenges for practical deployment.
+Addressing these computational burdens, while fully leveraging the semantic priors of SAM, is a primary motivation for our work.
 3 The Proposed Method
-3.1 Motivation and Problem Formulation
-Figure 4: Visual comparison of SOTA approaches across four commonly used datasets, including TNO, RoadScene, M3FD and FMB.
-The overall workflow of the proposed method is shown in Figure 3.
-Our objective is to develop a framework that effectively learns and leverages the semantic priors provided by SAM while minimizing the computational burden during inference.
-Integrating SAM’s semantic information can significantly enhance the quality of the fused images; however, directly involving SAM make the inference process computationally impractical due to its complexity.
-To solve this, we utilize a distillation scheme to transfer knowledge from a SAM-integrated main network to a lightweight sub-network.
-Traditional distillation approaches often face difficulties in capturing the intricate semantic relationships and maintaining the structural consistency necessary for tasks involving cross-modal information fusion.
-Considering these challenges and the need to jointly optimize both networks for effective semantic transfer, we propose a bi-level optimization framework that models the main and sub-networks as a unified system, aiming to enable a more cohesive and structured transfer of semantic insights.
-Denote that $\textbf{I}_{\text{vis}}$ and $\textbf{I}_{\text{ir}}$ represent the visible and infrared input images, respectively.
-Their corresponding semantic prior patches generated from SAM can be defined as $\textbf{P}_{\text{vis}}$ and $\textbf{P}_{\text{ir}}$.
-The main network, represented by $\mathcal{M}$ with parameters $\boldsymbol{\omega_{\mathrm{m}}}$.
-while the sub-network, represented by $\mathcal{S}$ with parameters $\boldsymbol{\omega_{\mathrm{s}}}$, aims to distill knowledge from the main network through the distillation process.
-Below our definition, the optimization process can be formulated as follows:
-$\min_{\omega_{\text{s}}}\mathcal{L}^{\text{d}}_{\text{m}}\big{(}\mathcal{M}(\{\textbf{I}_{\text{vis}},\textbf{I}_{\text{ir}}\};\{\textbf{P}_{\text{vis}},\textbf{P}_{\text{ir}}\};\boldsymbol{\omega_{\mathrm{m}}})\big{)},$ (1)
-$s.t.\boldsymbol{\omega_{\mathrm{s}}}^{*}\in\arg\min\limits_{\boldsymbol{\omega_{\mathrm{s}}}}\mathcal{L}^{\mathrm{d}}_{\text{s}}\big{(}\mathcal{S}(\mathbf{I_{\mathrm{vis}}},\mathbf{I_{\mathrm{ir}}};\boldsymbol{\omega_{\mathrm{s}}})\big{)},$ (2)
-where, $\mathcal{L}^{\mathrm{d}}_{\text{m}}$ represents the distillation loss for main network, ensuring high-quality fusion while embedding meaningful semantic information, and $\mathcal{L}^{\mathrm{d}}_{\text{s}}$ represents the distillation loss to train the sub-network for effectively mimicking the behavior of the main network.
-The optimized interaction process is reflected in fused refer image from $\mathcal{M}$, $\mathbf{I_{\mathrm{ref}}}=\mathcal{M}(\{\textbf{I}_{\text{vis}},\textbf{I}_{\text{ir}}\};\{\textbf{P}_{\text{vis}},\textbf{P}_{\text{ir}}\},\omega_{\text{m}})$ and the fused image $\mathbf{I_{\mathrm{f}}}=\mathcal{S}(\mathbf{I_{\mathrm{vis}}},\mathbf{I_{\mathrm{ir}}};\boldsymbol{\omega_{\mathrm{s}}}^{*})$.
-The goal is to jointly optimize both networks, where the main network learns the rich semantic priors, and the sub-network captures those features under reduced computational requirements.
-In the main network, source images are processed by SAM to generate image patches with semantic priors, which, along with features extracted from the source images at multiple scales, are passed through the SPA module to produce fused images.
-The sub-network serves as a distilled version of the main network.
-The bi-level optimization distillation scheme, involving triplet losses, facilitates the structured transfer of capabilities from the main network to the sub-network, ensuring that even with reduced complexity.
-The following sections will detail the design of the SPA module (Section 3.2) and the bi-level optimization distillation scheme involving triplet losses (Section 3.3).
+3.1 Problem Formulation
+Figure 4: Qualitative demonstrations of SOTA approaches across commonly used datasets, including TNO, RoadScene, M3FD and FMB.
+The overall workflow of our method is illustrated in Figure 3.
+Our aim is to fully exploit the semantic priors of SAM during the inference stage to enhance cross-modal fusion quality.
+However, directly deploying a large-scale SAM model often leads to prohibitive computational overhead.
+To mitigate this, we adopt a knowledge distillation strategy, transferring the information encoded by the SAM-driven main network to a lightweight sub-network, thereby significantly reducing inference costs while maintaining high-quality fusion.
+Nevertheless, the substantial capacity gap between the SAM-enhanced main network and the compact sub-network frequently results in incomplete semantic transfer or structural inconsistencies, which hinder ideal cross-modal fusion performance.
+To address this issue, we propose a bi-level optimization framework that jointly optimizes both networks as a unified system, aiming to bridge the distillation gap and maintain consistent fusion guided by SAM’s semantic priors.
+Let $\mathbf{I}_{\mathrm{s}}$ denote the source visible and infrared input images, i.e., $\mathbf{I}_{\mathrm{s}}=\{\mathbf{I}_{\mathrm{vis}},\mathbf{I}_{\mathrm{ir}}\}$.
+Their corresponding semantic-prior patches generated under the guidance of SAM are denoted as $\mathbf{P}_{\mathrm{s}}=\{\mathbf{P}_{\mathrm{vis}},\mathbf{P}_{\mathrm{ir}}\}$.
+Denote the main network by $\mathcal{M}$ with parameters $\boldsymbol{\omega_{\mathrm{m}}}$, and the sub-network by $\mathcal{S}$ with parameters $\boldsymbol{\omega_{\mathrm{s}}}$.
+We define the optimization process as:
+$\begin{gathered}\min_{\boldsymbol{\omega_{s}}}\mathcal{L}^{\mathrm{d}}_{\mathrm{s}}\big{(}\mathcal{S}(\mathbf{I}_{s},\,\boldsymbol{\omega_{s}});\boldsymbol{\omega_{m}^{*}}\big{)},\\ \text{s.t.}\quad\boldsymbol{\omega_{m}^{*}}\in\arg\min_{\boldsymbol{\omega_{m}}}\mathcal{L}^{\mathrm{d}}_{\mathrm{m}}\big{(}\mathcal{M}(\mathbf{I}_{s},\,\mathbf{P}_{s},\,\boldsymbol{\omega_{m}});\boldsymbol{\omega_{s}}\big{)},\end{gathered}$ (1)
+where $\mathcal{L}^{\mathrm{d}}_{\mathrm{m}}$ is the distillation loss that ensures high-quality fusion with meaningful semantic cues in the main network, and $\mathcal{L}^{\mathrm{d}}_{\mathrm{s}}$ is the distillation loss guiding the sub-network to effectively mimic the main network’s behavior.
+The fused reference image is given by: $\mathbf{I}_{\mathrm{ref}}=\mathcal{M}\big{(}\{\mathbf{I}_{\mathrm{vis}},\mathbf{I}_{\mathrm{ir}}\},\{\mathbf{P}_{\mathrm{vis}},\mathbf{P}_{\mathrm{ir}}\},\boldsymbol{\omega_{m}}\big{)}$, while the sub-network produces a fused image $\mathbf{I}_{\mathrm{f}}=\mathcal{S}\big{(}\mathbf{I}_{\mathrm{vis}},\mathbf{I}_{\mathrm{ir}},\boldsymbol{\omega_{s}}\big{)}$.
+Notably, both distillation objectives rely on outputs from the counterpart network, establishing a bidirectional dependency, which is a key feature of our distillation formulation.
+Within the overall framework, the main network integrates SAM-derived semantic priors into the fusion process, while the sub-network is optimized to align with its outputs under the bi-level formulation.
+The following sections elaborate on two key components that support this process: the Semantic Persistent Attention (SPA) module 3.2, which integrates SAM-derived priors with source features, and the triplet-based loss scheme 3.3, which facilitates structured knowledge transfer across networks.
 3.2 Semantic Persistent Attention Module
-To fully exploit the semantic information provided by SAM, we propose the Semantic Persistent Attention (SPA) Module, as illustrated in the flowchart in Figure. 3 (b).
+To fully exploit the semantic information provided by SAM, we propose the Semantic Persistent Attention (SPA) module, as illustrated in the flowchart in Figure 3 (b).
 The key core of the SPA module is a Persistent Repository (PR), which serves as a static memory to store and maintain crucial contextual information for the fusion process.
@@
 The core insight behind this design is that PR acts as a stable, modality-specific information source, guiding the cross-attention mechanism to fuse the semantic patches with rich contextual details from the original source.
-The visible and infrared semantic patches ($\textbf{P}_{\text{vis}}$ and $\textbf{P}_{\text{ir}}$) extracted from the Semantic Attention Module (SAM) are encoded into features ($\textbf{F}_{\text{pvis}}$ and $\textbf{F}_{\text{pir}}$).
+The visible and infrared semantic patches ($\textbf{P}_{\text{vis}}$ and $\textbf{P}_{\text{ir}}$) extracted from the SAM are encoded into features ($\textbf{F}_{\text{pvis}}$ and $\textbf{F}_{\text{pir}}$).
 These encoded semantic features represent limited portions of the scene and are processed through the cross-attention mechanism.
 By utilizing the key-value pairs ($\textbf{K}_{\text{src}},\textbf{V}_{\text{src}}$) stored in PR, the cross-attention mechanism enriches the semantic queries ($\textbf{Q}_{\text{pvis}}$ and $\textbf{Q}_{\text{pir}}$) with the full scene context from the source image, thereby addressing the inherent incomplete scene coverage in the patches.
-PR ensures that the fusion process remains grounded in a stable and consistent context, allowing the patches to be effectively supplemented with modality-specific information from the source.
-This stable context prevents the fusion process from drifting or losing semantic coherence, while still enabling flexibility in refining the features through attention mechanisms.
+PR ensures the fusion process stays grounded in a stable, consistent context, enabling patches to be effectively enriched with modality-specific information.
+This stability maintains semantic coherence while allowing flexible feature refinement via attention mechanisms.
 In summary, the SPA module uses PR to guide the cross-attention mechanism, ensuring that the semantic patches from SAM are enriched with consistent and detailed modality-specific information.
@@
 The main network heavily utilizes the SPA module, aiming to fully exploit the complex semantic priors provided by SAM, focus on capturing and retaining detailed semantic knowledge, thereby facilitating a powerful representation that can later be distilled into a more efficient sub-network.
-Figure 5: Quantitative comparison of fusion performance with other state-of-the-art methods on M3FD and FMB benchmarks.
-The Violin plots illustrating the distribution of metrics, in which the the black lines and white triangles indicate the medium and mean values.
+Figure 5: Quantitative comparison of fusion performance with other SOTA methods on M3FD and FMB benchmarks.
+The violin plots illustrate the distribution of metrics, in which the the black lines and white triangles indicate the medium and mean values.
 3.3 Triplet Loss in Distillation Scheme
-In this section, we introduce a triplet loss-driven distillation scheme that facilitates semantic knowledge transfer from the main network to the sub-network, where both networks are optimized in a bidirectional manner.
-This allows gradients to flow both from the main to the sub-network and vice versa, ensuring a cohesive transfer with mutual compromise.
-The distillation is driven by three distinct loss components, each targeting different aspects.
-The first loss operates at the feature level, aligning the reference image ($\mathbf{I}_{\text{ref}}$) and the fused image ($\mathbf{I}_{\text{f}}$) across multiple scales, which is defined as:
+In this section, we introduce a triplet loss-driven distillation scheme that facilitates semantic knowledge transfer from the main network ($\mathcal{M}$) to the sub-network ($\mathcal{S}$) under a bi-level optimization framework.
+Specifically, we adopt a DARTS-style [10] training protocol, where the main network (teacher) and the sub-network (student) are updated in small alternating steps, approximating the inner-outer structure of bilevel optimization.
+This allows gradients to flow in a bidirectional manner: the sub-network learns from the distillation signals provided by the main network, while the main network also adapts its parameters based on the sub-network’s performance and a segmentation objective.
+Consequently, the two networks reach a cohesive transfer with mutual compromise.
+Our distillation scheme is driven by three distinct losses, each focusing on a different aspect of the fused images.
+Let $\mathbf{I}_{\text{ref}}$ be the fused reference image produced by the main network, and $\mathbf{I}_{\text{f}}$ be the fused output from the sub-network.
+We first introduce a feature alignment term:
 $\mathcal{L}_{\text{fea}}=\sum_{m=1}^{M}\left(1-\frac{\textbf{F}_{\text{Den}}^{m}\cdot\textbf{F}_{\text{SPA}}^{m}}{\|\textbf{F}_{\text{Den}}^{m}\|_{2}\cdot\|\textbf{F}_{\text{SPA}}^{m}\|_{2}}\right),$ (2)
-where $\textbf{F}_{\text{Den}}^{m}$ and $\textbf{F}_{\text{SPA}}^{m}$ are the feature maps of the Dense and SPA blocks at the same scale $m$, and $\|\cdot\|_{2}$ is the l2-norm.
+where $\textbf{F}_{\text{Den}}^{m}$ and $\textbf{F}_{\text{SPA}}^{m}$ are the feature maps of the dense and SPA blocks at the same scale $m$, and $\|\cdot\|_{2}$ is the $\ell_{2}$-norm.
 The second loss component operates at the context level, which consists of two parts: a gradient part to preserve structural consistency and an MSE loss for intensity consistency. They are defined as:
 $\mathcal{L}_{\text{grad}}=\left\|\nabla\mathbf{I}_{\text{ref}}-\nabla\mathbf{I}_{\text{fus}}\right\|_{1},\quad\mathcal{L}_{\text{MSE}}=\left\|\mathbf{I}_{\text{ref}}-\mathbf{I}_{\text{fus}}\right\|_{2},$ (3)
-where $\nabla$ denotes the Sobel operator and $\|\cdot\|_{1}$ is the l1-norm. So the cont loss can be calculated as $\mathcal{L}_{\text{context}}=\mathcal{L}_{\text{grad}}+\mathcal{L}_{\text{MSE}}$.
+where $\nabla$ denotes the Sobel operator and $\|\cdot\|_{1}$ is the $\ell_{1}$norm.
+So the cont loss can be calculated as $\mathcal{L}_{\text{context}}=\mathcal{L}_{\text{grad}}+\mathcal{L}_{\text{MSE}}$.
+In practice, we apply these terms not only between $\mathbf{I}_{\text{ref}}$ and $\mathbf{I}_{\text{fus}}$, but also between each fused output and the original source images to ensure the reconstruction fidelity.
+This prevents both networks from drifting away from the source image during the distillation process.
 The third loss component, contrastive semantic loss aims to ensure that the feature space of the fused image remains close to that of the reference image, while being distinct from the individual visible and infrared images archiving efficient distillation.
@@
 The contrastive semantic loss on each modality is defined as:
-$\mathcal{L}_{\text{cs}}^{\text{ir}}=\sum_{l=1}^{L}\frac{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{fus}}\odot\mathbf{M}_{\text{ir}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ref}}\odot\mathbf{M}_{\text{ir}})\right\|_{1}^{2}}{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{x}}\odot\mathbf{M}_{\text{ir}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ir}}\odot\mathbf{M}_{\text{ir}})\right\|_{1}^{2}},$ (5)
-$\mathcal{L}_{\text{cs}}^{\text{vis}}=\sum_{l=1}^{L}\frac{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{fus}}\odot\mathbf{M}_{\text{vis}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ref}}\odot\mathbf{M}_{\text{vis}})\right\|_{1}^{2}}{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{x}}\odot\mathbf{M}_{\text{vis}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{vis}}\odot\mathbf{M}_{\text{vis}})\right\|_{1}^{2}},$ (6)
-where $l$ means $l$th layer of $\mathcal{S}_{\text{E}}$, and x $\in\{\text{ref},\text{fus}\}$.
+$\mathcal{L}_{\text{cs}}^{\text{ir}}=\sum_{l=1}^{L}\frac{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{fus}}\odot\mathbf{M}_{\text{ir}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ref}}\odot\mathbf{M}_{\text{ir}})\right\|_{2}}{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{x}}\odot\mathbf{M}_{\text{ir}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ir}}\odot\mathbf{M}_{\text{ir}})\right\|_{2}},$ (4)
+$\mathcal{L}_{\text{cs}}^{\text{vis}}=\sum_{l=1}^{L}\frac{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{fus}}\odot\mathbf{M}_{\text{vis}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ref}}\odot\mathbf{M}_{\text{vis}})\right\|_{2}}{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{x}}\odot\mathbf{M}_{\text{vis}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{vis}}\odot\mathbf{M}_{\text{vis}})\right\|_{2}},$ (5)
+where $l$ indexes the layers of $\mathcal{S}_{\text{E}}$, and x $\in\{\text{ref},\text{fus}\}$.
 So the total $\mathcal{L}_{\text{cs}}=\mathcal{L}_{\text{cs}}^{\text{ir}}+\mathcal{L}_{\text{cs}}^{\text{vis}}$ can also be computed bidirectionally in the bi-level optimization scheme between two networks, ensuring that the feature spaces of both the $\mathbf{I}_{\text{ref}}$ and the $\mathbf{I}_{\text{f}}$ are easily aligned.
 The total distillation loss for the sub-network, denoted as $\mathcal{L}^{\text{d}}_{\text{s}}$ in Eq.
-2, is the sum of the feature loss, context loss, and contrastive semantic loss: $\mathcal{L}^{\text{d}}_{\text{s}}=\mathcal{L}_{\text{fea}}+\mathcal{L}_{\text{cont}}+\mathcal{L}_{\text{cs}}$.
-For the main network, an additional segmentation cross-entropy loss $\mathcal{L}^{\text{seg}}$ is incorporated.
-This loss is computed between the segmentation predictions generated by an open-vocabulary SAM model $\mathcal{S}^{\text{O}}$ and the ground truth segmentation labels.
-By including this loss, we aim to leverage the capabilities of SAM to overcome the limitations of traditional segmentation networks, which typically rely on a fixed vocabulary.
+(1), is the sum of the feature loss, context loss, and contrastive semantic loss: $\mathcal{L}^{\text{d}}_{\text{s}}=\mathcal{L}_{\text{fea}}+\mathcal{L}_{\text{cont}}+\mathcal{L}_{\text{cs}}$.
+For the main network, an additional segmentation cross-entropy loss $\mathcal{L}_{\text{seg}}$ is incorporated.
+This loss is computed between the segmentation predictions generated by an open-vocabulary model $\mathcal{S}_{\text{O}}$ [45, 46] and the ground truth segmentation labels.
 Instead of affecting the sub-network, this loss aim to prevent potential optimization conflicts.
-The segmentation loss $\mathcal{L}^{\text{seg}}$ is defined as:
-$\mathcal{L}^{\text{seg}}=-\sum_{c}\left[\mathbf{Y}_{\text{label}}^{c}\log(\mathbf{Y}_{\text{pred}}^{c})\right],$ (7)
+The segmentation loss $\mathcal{L}_{\text{seg}}$ is defined as:
+$\mathcal{L}_{\text{seg}}=-\sum_{c}\left[\mathbf{Y}_{\text{label}}^{c}\log(\mathbf{Y}_{\text{pred}}^{c})\right],$ (6)
 where $\mathbf{Y}_{\text{label}}$ represents the label segmentation map, $\mathbf{Y}_{\text{pred}}$ is the predicted segmentation map, and $c$ is the class index.
-For now, the total distillation loss for the main network can be calculated as: $\mathcal{L}^{\text{d}}_{\text{m}}=\mathcal{L}^{\text{d}}_{\text{s}}+\mathcal{L}^{\text{seg}}$.
+For now, the total distillation loss for the main network can be calculated as: $\mathcal{L}^{\text{d}}_{\text{m}}=\mathcal{L}^{\text{d}}_{\text{s}}+\mathcal{L}_{\text{seg}}$.
 4 Experiments
@@
 SAGE Proposed 89.0 83.9 71.4 86.3 66.7 64.1 81.7 61.2 81.0 70.4 17.0 75.2 77.7 64.0 82.8 51.1
-Table 1: Quantitative semantic segmentation IoU results of different methods on the FMB dataset.
-Figure 6: Visual comparison of segmentation results of different methods on the FMB dataset.
+Table 1: Quantitative semantic segmentation results of SOTA approaches on the FMB dataset.
+Figure 6: Qualitative segmentation results of SOTA approaches on the FMB dataset.
 4.1 Settings and Details
 Five representative datasets, namely TNO [27], RoadScene [32], MFNet [2], FMB [13], and M3FD [11], are utilized for both training and evaluation.
-For fusion-based segmentation evaluations, SegFormer [41] (B3 variant) is adopted as the backbone, and the models are trained for 100 epochs.
+For segmentation, SegFormer [31] (B3 variant) is adopted as the backbone, and the models are trained for 100 epochs.
 The training and testing splits are strictly in accordance with the official dataset guidelines.
-The Adam optimizer is used to optimize both networks during the training period, with learning rates of $5\times 10^{-5}$ for both networks.
+Adam is adopted for training, with initial learning rates of $5\times 10^{-4}$ and $2\times 10^{-3}$ for the main and sub-networks, respectively.
+Cosine annealing decays both to $1\times 10^{-5}$.
 Prior to distillation, the models undergo a pre-training phase, followed by 5 epochs of distillation.
-The batch size is set to 4, and the images are randomly cropped and resized to $192\times 160$ during training.
+The batch size is set to 4, and the images are randomly cropped and resized to $192\times 256$ during training.
 The entire framework is implemented in PyTorch and executed on two NVIDIA GeForce RTX 4090 GPUs.
-4.2 Results of IVIF
+4.2 Results of Multi-modality Image Fusion
 We demonstrate our fusion quality by visualization and quantitative analysis with 9 other SOTA methods in recent years, including DDFM [40], U2Fusion [32], TarDAL [11], SegMiF [13], FILM [42], SHIP [43], MRFS [36], EMMA [41] and TIMFusion [19].
-Visualization Comparisons.
-Figure.
-4 presents a visual comparison of various methods.
+Qualitative Comparisons.
+Figure 4 presents a visual comparison of various methods.
 Overall, the proposed method has two main advantages.
@@
 The integration of SAM enhances our method, giving it the ability to surpass other methods and achieve superior results.
-Quantitative Results.
-We also compare our numerical results with other competitors in Figure.
-5, based on 100 image pairs randomly selected from the FMB and M3FD datasets.
-We employ four fusion evaluation metrics, including the Entropy (EN) [31], Standard Deviation (SD) [1], Sum of Correlation Differences (SCD) [24], and Multi-Scale Structural Similarity (MS-SSIM) [39].
-Additionally, we introduce four No-Reference Image Quality Assessments (NR-IQA) related to image quality: BRISQUE [27], NIQE [28], MUSIQ [7], and PaQ-2-PiQ [43].
+Quantitative Comparisons.
+We also compare our numerical results with other competitors in Figure 5, based on 100 image pairs randomly selected from the FMB and M3FD datasets.
+We employ four fusion evaluation metrics [21], including the Entropy (EN), Standard Deviation (SD), Sum of Correlation Differences (SCD), and Multi-Scale Structural Similarity (MS-SSIM).
+Additionally, we introduce four no-reference image quality assessments [34] related to image quality: BRISQUE, NIQE, MUSIQ, and PaQ-2-PiQ.
 In terms of these metrics, our method demonstrates consistent superiority.
@@
 Framework Segformer - B3 [31]
-Method Venue Car Per. Bike Curve C.S. C.C. Bump mIoU
-Visible - 83.3 54.7 59.6 7.46 31.0 37.4 30.8 44.6
-Infrared - 79.0 63.8 51.4 7.10 25.2 34.1 41.8 44.4
-DDFM ICCV’23 86.7 62.2 63.4 30.8 35.5 45.7 41.1 51.7
-U2Fusion TPAMI’22 87.3 65.9 60.8 33.4 36.0 46.1 40.9 52.0
-TarDAL CVPR’22 86.4 67.1 62.4 28.4 30.5 45.0 41.5 51.0
-SegMiF ICCV’23 85.9 66.3 64.4 34.0 34.7 43.0 48.0 52.4
-FILM ICML’24 86.6 66.4 61.9 36.3 32.0 46.7 46.7 52.7
-SHIP CVPR’24 86.7 63.9 59.7 33.5 35.2 44.4 42.4 51.5
-MRFS CVPR’24 86.1 61.5 59.7 30.6 35.2 44.9 38.4 50.4
-EMMA CVPR’24 86.1 65.6 63.8 25.4 29.1 44.9 41.3 51.7
-TIMFusion TPAMI’24 86.4 58.7 60.3 29.8 33.6 44.1 38.3 49.8
-SAGE Proposed 87.9 67.5 64.3 41.2 34.1 46.8 46.7 54.0
-Table 2: Quantitative semantic segmentation results of different methods on the MFNet dataset.
-Figure 7: Visual comparison of segmentation results of different methods on the MFNet dataset.
-4.3 Results of IVIF based Segmentation
-Visualization Comparisons.
-We present the segmentation visualization results on the challenging new FMB dataset in Figure.
-6.
+Method Car Per. Bike Curve C.S. C.C. Bump mIoU
+Visible 83.3 54.7 59.6 7.46 31.0 37.4 30.8 44.6
+Infrared 79.0 63.8 51.4 7.10 25.2 34.1 41.8 44.4
+DDFM 86.7 62.2 63.4 30.8 35.5 45.7 41.1 51.7
+U2Fusion 87.3 65.9 60.8 33.4 36.0 46.1 40.9 52.0
+TarDAL 86.4 67.1 62.4 28.4 30.5 45.0 41.5 51.0
+SegMiF 85.9 66.3 64.4 34.0 34.7 43.0 48.0 52.4
+FILM 86.6 66.4 61.9 36.3 32.0 46.7 46.7 52.7
+SHIP 86.7 63.9 59.7 33.5 35.2 44.4 42.4 51.5
+MRFS 86.1 61.5 59.7 30.6 35.2 44.9 38.4 50.4
+EMMA 86.1 65.6 63.8 25.4 29.1 44.9 41.3 51.7
+TIMFusion 86.4 58.7 60.3 29.8 33.6 44.1 38.3 49.8
+SAGE 87.9 67.5 64.3 41.2 34.1 46.8 46.7 54.0
+Table 2: Quantitative semantic segmentation results of SOTA approaches on the MFNet dataset.
+Figure 7: Qualitative segmentation results of SOTA approaches on the MFNet dataset.
+4.3 Results of Multi-modality Segmentation
+Qualitative Comparisons.
+We present the segmentation visualization results on the challenging new FMB dataset in Figure 6.
 In the daytime intersection scene, thanks to the powerful semantic priors provided by SAM, our method is the only one to completely distinguish between the truck and the bus.
 In the nighttime road scene, we successfully segment the sidewalk, achieving optimal performance.
-Additionally, we showcase the segmentation results on the MFNet dataset in Figure.
-7.
+Additionally, we showcase the segmentation results on the MFNet dataset in Figure 7.
 Leveraging the advanced semantic information learned by SAM, our method successfully segments distant small pedestrian targets during the daytime and near-invisible lane curve at night.
-Quantitative Results.
-Table.
-1 presents a qualitative comparison of IoU results for two segmentation frameworks on the FMB dataset.
-The first framework involves re-training images generated by various fusion methods, while the second uses the output from the open-vocabulary segmentation network with prompt words, without requiring retraining.
-In traditional comparisons, our method achieves a 3.0 mIoU improvement over the second-best approach, demonstrating its competitive edge across multiple categories.
-Additionally, in the segmentation network that requires no training, our method also performs favorably, benefiting from the efficient adaptability provided by SAM’s semantic information.
-Table.
-2 further showcases the qualitative results on the MFNet dataset, which has lower resolution and fewer labels.
-Although our method shows some competitive gaps in background categories (e.g., bicycles, colored cones), it achieves significant overall superiority in key perception categories (e.g., Car and Person) and mean performance.
-Our method seamlessly integrates SAM with the fusion network, achieving a breakthrough in performance for efficient segmentation task.
-Variants mIoU Metric
+Quantitative Comparisons.
+Table 1 presents a qualitative comparison of IoU results for two segmentation frameworks on the FMB dataset.
+The first framework involves re-training images generated by various fusion methods, while the second uses the output from an open-vocabulary segmentation network with prompt words, without retraining.
+In traditional comparisons, our method achieves a 3.0 mIoU improvement over the second-best approach, demonstrating its competitive edge across categories.
+Additionally, in the segmentation network that requires no training, our method also performs favorably, benefiting from the adaptability provided by SAM’s semantic information.
+Table 2 further showcases the qualitative results on the MFNet dataset, which has lower resolution and fewer labels.
+Although our method shows some gaps in background categories, it achieves overall superiority in key perception categories and mean performance.
+Our method integrates SAM with the fusion network, achieving a breakthrough in performance for efficient segmentation.
+Variants mIoU Fusion Metrics
 MFNet / FMB EN / SD/ SCD / MS-SSIM
-I. SAM in the Pipeline
-(a) w/o SAM in $\mathcal{M}$ 50.7 / 56.5 6.713 / 36.93 / 1.612 / 1.127
+I. The Impact of SAM
+(a) w/o SAM 50.7 / 56.5 6.713 / 36.93 / 1.612 / 1.127
 (b) P from $\mathbf{Y}_{\text{label}}$ 53.2 / 56.4 6.239 / 33.52 / 1.412 / 1.102
 (c) Swap $\mathcal{S}^{\text{O}}$ 53.4 / 55.5 6.312 / 34.79 / 1.316 / 1.097
-II. Design of SPA
+II. Study on the SPA Module
 (a) w/o Z 52.3 / 57.2 6.333 / 32.81 / 1.384 / 1.002
@@
 (c) w/o PR 53.3 / 57.2 6.556 / 35.48 / 1.576 / 1.046
-III. Distillation Triplet Losses
+III. Discussion on the Distillation Scheme
 (a) w/o $\mathcal{L}_{\text{fea}}$ 49.7 / 51.6 6.436 / 37.28 / 1.277 / 1.079
@@
 (c) w/o $\mathcal{L}_{\text{cs}}$ 51.7 / 54.3 6.524 / 39.11 / 1.356 / 1.038
-(d) Online Dist. 50.0 / 50.4 6.221 / 32.65 / 1.581 / 0.987
-The proposed method 54.0 / 61.2 6.872 / 40.77 / 1.604 / 1.117
+(d) Offline Dist. 50.0 / 50.4 6.221 / 32.65 / 1.581 / 0.987
+SAGE 54.0 / 61.2 6.872 / 40.77 / 1.604 / 1.117
 Table 3: Quantitative results on both fusion and segmentation of all variants in three ablation studies.
@@
 4.4 Ablation Studies
-Impact of SAM.
+The Impact of SAM.
 The Segment Anything Module serves as a critical cornerstone in our method.
@@
 In variant (b), we assist the generation of semantic patches with segmentation labels.
-Additionally, we replace $\mathcal{S}^{\text{O}}$ with a conventional segmentation network [41].
-Variant (a) enhances the source image information, leading to the optimal similarity metrics, SCD and MS-SSIM, as shown in Table.
-3.
-Our method fully leverages the semantic priors provided by SAM, resulting in the best foggy building contours, as illustrated in the first row of Figure.
-8.
+Additionally, we replace $\mathcal{S}_{\text{O}}$ with a conventional segmentation network [31].
+Variant (a) enhances the source image information, leading to the optimal similarity metrics, SCD and MS-SSIM, as shown in Table 3.
+Our method fully leverages the semantic priors provided by SAM, resulting in the best foggy building contours, as illustrated in the first row of Figure 8.
 Figure 8: Visualization comparison of impact of SAM.
 Figure 9: Visualization comparison of ablation on SPA.
-Study on SPA module.
+Study on the SPA Module.
 The proposed Semantic Persistent Attention module plays a key role in integrating the semantic priors from SAM, effectively guiding the retention and enhancement of inherent modality features.
-To investigate its impact, we set up three variants: (a) w/o latent representation (Z), (b) w/o key-value pairs in the Persistent Repository (PR), and (c) w/o PR.
+To investigate its impact, we set up three variants: (a) w/o latent representation Z, (b) w/o key-value pairs in the Persistent Repository, and (c) w/o PR.
 The channel loss due to these modifications has been corrected.
-A visual comparison of these ablation variants is shown in Figure.
-9.
+A visual comparison of these ablation variants is shown in Figure 9.
 It is evident that PR in the SPA module plays an essential role in maintaining critical information.
@@
 Similarly, the absence of PR in the SPA module completely prevents the capture of beneficial semantic information, causing the network to lose focus on key areas and resulting in a low-contrast fusion results in (c).
-Discussion on distillation scheme.
+Figure 10: Visualization comparison of ablation on Distillation.
+Discussion on the Distillation Scheme.
 Furthermore, we discuss the distillation scheme.
-We first conduct ablation studies on each component of the triple loss function, yielding variants (a)-(c) in Table.
-3 III.
-We also replace the dual-layer optimization distillation with a traditional online distillation approach, forming variant (d).
-A visual comparison is shown in Figure.
-10.
-Notably, traditional online distillation results are significantly inferior to our method, both in visual quality and the gradient map from variant (b).
-This confirms the effectiveness of our distillation method when incorporating semantic information.
-Figure 10: Visualization comparison of ablation on Distillation.
+First, we conduct ablation studies on each component of the triple loss function, leading to variants (a)-(c) as shown in Table 3 III.
+Additionally, we replace the bi-level optimization distillation method with the offline distillation approach, forming variant (d).
+A visual comparison of these variants is shown in Figure 10.
+Notably, the results of the traditional online distillation are significantly inferior to our method, both in terms of visual quality and the gradient map generated by variant (b).
+This confirms the effectiveness of the proposed distillation method when incorporating semantic information.
 4.5 Computational Efficiency Analysis
 In the M3FD benchmark, we compare our method with 9 other SOTA methods in terms of time, FLOPS, and parameters.
-As shown in Table.
-4, our method demonstrates significant advantages across all aspects, particularly in time and FLOPS.
+As shown in Table 4, our method demonstrates significant advantages across all aspects, particularly in time and FLOPs.
 Specifically, due to our distillation scheme, our sub-network is able to adjust flexibly while maintaining high efficiency, resulting in reduced FLOPs.
@@
 Extensive experiments show that our method achieves superior fusion and task performance, outperforming across multiple datasets.
+Acknowledgments
+This work is partially supported by the National Natural Science Foundation of China (Nos. 62450072, U22B2052,
+62302078), Central Guidance for Local Science and Technology Development Fund (Youth Science Fund Project, Category A, No. 2025JH6/101100001), the Distinguished
+Young Scholars Funds of Dalian (No. 2024RJ002), the Chi-
+na Postdoctoral Science Foundation (No. 2023M730741) and the Fundamental Research Funds for the Central Universities.